If sam can do a job in 4 days that Lisa can do in 6 days, and Tom can do in 2 days, how long would the job take if Sam, Lisa, and Tom work together to complete it?
Answer by nerdybill(7384)

If sam can do a job in 4 days that Lisa can do in 6 days, and Tom can do in 2 days, how long would the job take if Sam, Lisa, and Tom work together to complete it?
.
Let x = time (days) it takes for all three
then
x(1/4 + 1/6 + 1/2) = 1
multiplying both sides by 12:
x(3 + 2 + 6) = 12
x(11) = 12
x = 12/11
x = 1.09 days
or
x = 1 day, 2 hours and 11 minutes